OpenStudy (anonymous):

I just recieved this question on homework and have been working on it for quite some. Need help! Evaluate the following expression using the values given: Find 3x2 - y3 - y3 - z if x = 3, y = -2, and z = -5. Numerical Answers Expected!

OpenStudy (anonymous):

\[3x^2 - y^3 - y^3 - z\] if \[x = 3, y = -2, and z = -5.\]

OpenStudy (igreen):

@satellite73

OpenStudy (igreen):

OS is messed up, can you fix it?

OpenStudy (igreen):

@Preetha

OpenStudy (anonymous):

no

OpenStudy (igreen):

What's wrong with it?

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

first off \[3x^2 - y^3 - y^3 - z=3x^2-2y^3-z\]

OpenStudy (anonymous):

Im sorry im a little confused.

OpenStudy (anonymous):

then \[3\times 3^2-2\times (-2)^3-(-5)\] is what you need to compute

OpenStudy (anonymous):

Ah I see

OpenStudy (anonymous):

i would use a calculator http://www.wolframalpha.com/input/?i=3*3^2-2*+%28-2%29^3-%28-5%29

OpenStudy (anonymous):

So 3x3^2 - 2 x (-2)^3 - (-5) Should give me my answer?

OpenStudy (anonymous):

So my answer is 48?

OpenStudy (anonymous):

So my answer would be 48?

OpenStudy (anonymous):

@satellite73

OpenStudy (texaschic101):

3x^2 - y^3 - y^3 - z when x = 3, y = -2 and z = -5 lets sub 3(3^2) - (-2^3) - (-2^3) -(-5) 3(9) - (-8) - (-8) + 5 27 + 8 + 8 + 5 27 + 21 = 48 thats what I got

OpenStudy (anonymous):

Oh I see thank you! I got the same answer.

OpenStudy (texaschic101):

great minds think alike :)

OpenStudy (anonymous):

Lol.